# Channel Knowledge Map Construction via Physics-Inspired Diffusion Model Without Prior Observations

### v1

@@ IV-C The Physics-inspired Channel Knowledge Map Generation Diffusion Model Architecture and Algorithm Flow
V Experiment Results
V-A Experiment Setup
V-A1 Dataset
In this paper, the open-source dataset named RadioMapSeer[47] is utilized. The dataset was collected from 700 urban maps in European cities. For each city map, a single base station was deployed, and Winprop was employed to perform ray-tracing channel simulations and generate the corresponding CGM heatmaps. Each CGM is discretized into a $256\times 256$ grid, where each grid cell represents the path loss value at the corresponding location. RadioMapSeer provides three groups of datasets obtained by different simulation methods, namely DPM (Dominant Path Model), IRT2 (Intelligent Ray Tracing with 2 interactions), and IRT4 (Intelligent Ray Tracing with 4 interactions). For each method, both static CGMs (without vehicles) and dynamic CGMs (with vehicles) were simulated. In the DPM-based dataset, 80 different base station locations were randomly selected for each urban map, resulting in a total of $80\times 700=56000$ CGMs. In our experiments, we use the DPM-generated dataset, which is divided into 40,000 CGMs for training, 8,000 CGMs for validation, and 8,000 CGMs for testing.
V-A2 Implentation Details
@@ V-A3 Evaluation Metric
Figure 8: Comparison of Static CGMs Constructed by Different Models.
V-A4 Benchmark
@@
Both RadioDiff and the proposed model are based on the diffusion model framework. However, unlike our model, RadioDiff adopts the training and generation algorithms from the Decoupled Diffusion Model (DDM)[48]. This framework requires a neural network with a dual UNet architecture. Moreover, to enhance the sharpness of the reconstructed CGM edges, an Adaptive Fast Fourier Transform (FFT) module is introduced between the encoder and decoder of the UNet. As a result, the number of network parameters in RadioDiff is significantly higher than that of the proposed model.
@@
V-B Performance Comparison
V-B1 Static Channel Gain Map
The evaluation metrics of static CGM reconstruction on the test set for different models are presented in Table I. It can be observed that the proposed model achieves the best performance across all evaluation metrics. The reconstructed CGMs are visualized in grayscale images, where darker regions indicate higher path loss values. We compare the CGMs reconstructed by different models with the ground-truth CGMs, as shown in Fig. (8), where the black regions represent buildings.
TABLE I: Quantitative comparison of static CGM reconstruction results among different models
@@ V-B1 Static Channel Gain Map
As shown in Fig. (8), the CGMs reconstructed by RadioUNet, a discriminative model, exhibit very blurry edges and fail to capture the shadowing effect caused by obstacles in wireless signal propagation. The RME-GAN, which is a generative model based on GANs, produces CGMs with sharper edges; however, due to the inherent training instability of GAN and their limited capability in fitting complex data distributions, the reconstructed shadowed regions and edges deviate significantly from the ground-truth CGMs. Consequently, RME-GAN yields the worst performance among all compared models in terms of evaluation metrics. The RadioDiff, leveraging the strong generative capability of diffusion models, alleviates the blurry-edge problem observed in RadioUNet and improves the accuracy of shadowed regions compared with RME-GAN. However, since RadioDiff employs a model originally designed for the image domain, the constructed CGM, though visually plausible, lacks physical realism. In local regions, the edges of occluded areas remain distorted, and the spatial distribution of regions heavily blocked by obstacles is still physically unreasonable. In contrast, the CGMs reconstructed by the proposed physics-informed model exhibit spatial distributions that are more consistent with the underlying propagation environment. The grayscale details of the reconstructed CGMs better reflect the physical characteristics of wireless signal attenuation, thereby achieving the highest reconstruction accuracy among all models.
V-B2 Dynamic Channel Gain Map
Table II presents the quantitative results of dynamic CGM reconstruction for different models. It can be observed that the proposed model also achieves the best performance in constructing dynamic CGMs, with a larger performance gain compared to the static case. This is because, in dynamic CGMs, vehicles in addition to buildings act as obstacles, making the propagation environment more complex. By introducing physical constraints, the proposed model ensures stronger physical consistency between the generated dynamic CGMs and the actual propagation environment, thereby achieving higher reconstruction accuracy.
TABLE II: Quantitative comparison of dynamic CGM reconstruction results among different models
@@ V-B2 Dynamic Channel Gain Map
Fig. (9) shows the comparison between the dynamic CGMs reconstructed by different models and the ground-truth dynamic CGMs, where the blue dotted regions represent vehicles. As observed from Fig. (9), the dynamic CGMs reconstructed by RadioUNet still exhibit blurry edges and fail to effectively capture the partial shadowing effect caused by vehicles. Although RME-GAN produces relatively sharper edges at the locations of vehicles, it introduces large reconstruction errors in the spatial distribution of shadowed regions. RadioDiff is able to accurately reflect the shadowed areas caused by vehicles but still lacks fine-grained details. In contrast, the proposed physics-informed model, benefiting from the introduced physical constraints, can more precisely capture the subtle path loss variations caused by small obstacles such as vehicles.
V-B3 Parameter Size and Construction Time
TABLE III: Comparison of network parameter scales and reconstruction times among different models
@@ V-B3 Parameter Size and Construction Time
Table III presents the comparison of network parameter scales and reconstruction times among the four CGM construction models. Since diffusion-based models involve additional components such as a VAE, a conditional encoder, and cross-attention mechanisms, they generally have larger network sizes and longer inference times. However, the proposed model achieves the highest reconstruction accuracy while maintaining a smaller number of parameters and shorter reconstruction time compared with RadioDiff. Moreover, considering that variations in CGM are mainly caused by the movement of vehicles or pedestrians, the temporal interval of such changes is typically no less than 1s. With an average reconstruction time of approximately 0.68s, the proposed model is capable of satisfying the real-time CGM construction requirement and providing high-precision CGMs for downstream communication tasks.
V-C Ablation study
@@
TABLE IV: Quantitative comparison of construction results with different physical loss terms introduced
@@
In this section, we compare the experimental results obtained under four configurations: without any physics-inspired terms, with one, with two, and with all three physics-inspired terms, as shown in Table IV. For simplicity, the edge loss, region propagation loss, and multi-scale feature loss are denoted as $L_{1}$, $L_{2}$ and $L_{3}$, respectively. ”None” represents the case where no physics-inspired terms are introduced, i.e., the vanilla LDM is used for CGM construction, which yields the highest reconstruction error according to Table IV.
@@ VI Conclusion
[Derivation of the Equation (29)] The detailed derivation of Equation 29 is given as follows:
$E_{\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right)}\left[-\log p_{\theta}\left(\mathbf{x}_{0}\right)\right]$
$+E_{\mathbf{x}_{1:T}\sim q\left(\mathbf{x}_{1:T}\right)}\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]$
$\leq$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]$
$+E_{\mathbf{x}_{1:T}\sim q\left(\mathbf{x}_{1:T}\right)}\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]$
$=$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]$
$+\int\cdots\int\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]q\left(\mathbf{x}_{1:T}\right)d\mathbf{x}_{1:T}$
$=$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]$
$+\int\cdots\int\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]\int q\left(\mathbf{x}_{0:T}\right)d\mathbf{x}_{0}d\mathbf{x}_{1:T}$
$=$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]$
$+\int\cdots\int\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]q\left(\mathbf{x}_{0:T}\right)d\mathbf{x}_{0:T}$
@@
$=$ $E_{t\sim[1,T],\mathbf{x}_{0:T}\sim q\left(\mathbf{x}_{0:T}\right)}[\frac{\tilde{A}_{t}^{2}}{2\tilde{\beta}_{t}^{2}}\left\|\mathbf{x}_{0}-\mathbf{x}_{\theta}\left(\mathbf{x}_{t},t\right)\right\|^{2}$
$+\frac{1}{2\sigma^{2}}\left\|r\left(\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{t},t\right)\right)\right\|^{2}],$
where, the detailed proof of the following equation can be found in [43].
$E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]$
$=$ $E_{t\sim[1,T],\mathbf{x}_{0:T}\sim q\left(\mathbf{x}_{0:T}\right)}\left[\frac{\tilde{A}_{t}^{2}}{2\tilde{\beta}_{t}^{2}}\left\|\mathbf{x}_{0}-\mathbf{x}_{\theta}\left(\mathbf{x}_{t},t\right)\right\|^{2}\right].$

### v2

@@ 4.3 The Physics-inspired Channel Knowledge Map Generation Diffusion Model Architecture and Algorithm Flow
5 Experiment Results
5.1 Experiment Setup
5.1.1 Dataset
In this paper, the open-source dataset named RadioMapSeer[0gtx6v3022] is utilized. The dataset was collected from 700 urban maps in European cities. For each city map, a single BS was deployed, and Winprop was employed to perform ray-tracing channel simulations and generate the corresponding CGM heatmaps. Each CGM is discretized into a $256\times 256$ grid, where each grid cell represents the path loss value at the corresponding location. RadioMapSeer provides three groups of datasets obtained by different simulation methods, namely DPM (Dominant Path Model), IRT2 (Intelligent Ray Tracing with 2 interactions), and IRT4 (Intelligent Ray Tracing with 4 interactions). For each method, both static CGMs (without vehicles) and dynamic CGMs (with vehicles) were simulated. In the DPM-based dataset, 80 different BS locations were randomly selected for each urban map, resulting in a total of $80\times 700=56000$ CGMs. In our experiments, we use the DPM-generated dataset, which is divided into 40,000 CGMs for training, 8,000 CGMs for validation, and 8,000 CGMs for testing.
5.1.2 Implentation Details
@@ 5.1.3 Evaluation Metric
5.1.4 Benchmark
@@
Both RadioDiff and the proposed model are based on the diffusion model framework. However, unlike our model, RadioDiff adopts the training and generation algorithms from the Decoupled Diffusion Model (DDM)[huang2023decoupled]. This framework requires a neural network with a dual UNet architecture. Moreover, to enhance the sharpness of the reconstructed CGM edges, an Adaptive Fast Fourier Transform (FFT) module is introduced between the encoder and decoder of the UNet. As a result, the number of network parameters in RadioDiff is significantly higher than that of the proposed model.
@@
5.2 Performance Comparison
5.2.1 Static Channel Gain Map
The evaluation metrics of static CGM reconstruction on the test set for different models are presented in Table 1. It can be observed that the proposed model achieves the best performance across all evaluation metrics. The reconstructed CGMs are visualized in grayscale images, where darker regions indicate higher path loss values. We compare the CGMs reconstructed by different models with the ground-truth CGMs, as shown in Fig. (8), where the black regions represent buildings.
Figure 8: Comparison of Static CGMs Constructed by Different Models.
Table 1: Quantitative comparison of static CGM reconstruction results among different models
@@ 5.2.1 Static Channel Gain Map
As shown in Fig. (8), the CGMs reconstructed by RadioUNet, a discriminative model, exhibit very blurry edges and fail to capture the shadowing effect caused by obstacles in wireless signal propagation. The RME-GAN, which is a generative model based on GANs, produces CGMs with sharper edges; however, due to the inherent training instability of GAN and their limited capability in fitting complex data distributions, the reconstructed shadowed regions and edges deviate significantly from the ground-truth CGMs. Consequently, RME-GAN yields the worst performance among all compared models in terms of evaluation metrics. The RadioDiff, leveraging the strong generative capability of diffusion models, alleviates the blurry-edge problem observed in RadioUNet and improves the accuracy of shadowed regions compared with RME-GAN. However, since RadioDiff employs a model originally designed for the image domain, the constructed CGM, though visually plausible, lacks physical realism. In local regions, the edges of occluded areas remain distorted, and the spatial distribution of regions heavily blocked by obstacles is still physically unreasonable. In contrast, the CGMs reconstructed by the proposed physics-inspired model exhibit spatial distributions that are more consistent with the underlying propagation environment. The grayscale details of the reconstructed CGMs better reflect the physical characteristics of wireless signal attenuation, thereby achieving the highest reconstruction accuracy among all models.
5.2.2 Dynamic Channel Gain Map
Table 2 presents the quantitative results of dynamic CGM reconstruction for different models. It can be observed that the proposed model also achieves the best performance in constructing dynamic CGMs, with a larger performance gain compared to the static case. This is because, in dynamic CGMs, vehicles in addition to buildings act as obstacles, making the propagation environment more complex. By introducing physical constraints, the proposed model ensures stronger physical consistency between the generated dynamic CGMs and the actual propagation environment, thereby achieving higher reconstruction accuracy.
Table 2: Quantitative comparison of dynamic CGM reconstruction results among different models
@@ 5.2.2 Dynamic Channel Gain Map
Fig. (9) shows the comparison between the dynamic CGMs reconstructed by different models and the ground-truth dynamic CGMs, where the blue dotted regions represent vehicles. As observed from Fig. (9), the dynamic CGMs reconstructed by RadioUNet still exhibit blurry edges and fail to effectively capture the partial shadowing effect caused by vehicles. Although RME-GAN produces relatively sharper edges at the locations of vehicles, it introduces large reconstruction errors in the spatial distribution of shadowed regions. RadioDiff is able to accurately reflect the shadowed areas caused by vehicles but still lacks fine-grained details. In contrast, the proposed physics-inspired model, benefiting from the introduced physical constraints, can more precisely capture the subtle path loss variations caused by small obstacles such as vehicles.
5.2.3 Parameter Size and Construction Time
Table 3: Comparison of network parameter scales and reconstruction times among different models
@@ 5.2.3 Parameter Size and Construction Time
Table 3 presents the comparison of network parameter scales and reconstruction times among the four CGM construction models. Since diffusion-based models involve additional components such as a VAE, a conditional encoder, and cross-attention mechanisms, they generally have larger network sizes and longer inference times. However, the proposed model achieves the highest reconstruction accuracy while maintaining a smaller number of parameters and shorter reconstruction time compared with RadioDiff. Moreover, considering that variations in CGM are mainly caused by the movement of vehicles or pedestrians, the temporal interval of such changes is typically no less than 1s. With an average reconstruction time of approximately 0.68s, the proposed model is capable of satisfying the real-time CGM construction requirement and providing high-precision CGMs for downstream communication tasks.
5.3 Ablation study
@@
In this section, we compare the experimental results obtained under four configurations: without any physics-inspired terms, with one, with two, and with all three physics-inspired terms, as shown in Table 4. For simplicity, the edge loss, region propagation loss, and multi-scale feature loss are denoted as $L_{1}$, $L_{2}$ and $L_{3}$, respectively. ”None” represents the case where no physics-inspired terms are introduced, i.e., the vanilla LDM is used for CGM construction, which yields the highest reconstruction error according to Table 4.
Table 4: Quantitative comparison of construction results with different physical loss terms introduced
@@ 6 Conclusion
7 Appendix
The detailed derivation of Equation 29 is given as follows:
$E_{\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right)}\left[-\log p_{\theta}\left(\mathbf{x}_{0}\right)\right]+E_{\mathbf{x}_{1:T}\sim q\left(\mathbf{x}_{1:T}\right)}\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]$
$\leq$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]+E_{\mathbf{x}_{1:T}\sim q\left(\mathbf{x}_{1:T}\right)}\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]$
$=$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]+\int\cdots\int\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]q\left(\mathbf{x}_{1:T}\right)d\mathbf{x}_{1:T}$
$=$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]+\int\cdots\int\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]\int q\left(\mathbf{x}_{0:T}\right)d\mathbf{x}_{0}d\mathbf{x}_{1:T}$
$=$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]+\int\cdots\int\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]q\left(\mathbf{x}_{0:T}\right)d\mathbf{x}_{0:T}$
@@ 7 Appendix
$=$ $E_{t\sim[1,T],\mathbf{x}_{0:T}\sim q\left(\mathbf{x}_{0:T}\right)}[\frac{\tilde{A}_{t}^{2}}{2\tilde{\beta}_{t}^{2}}\left\|\mathbf{x}_{0}-\mathbf{x}_{\theta}\left(\mathbf{x}_{t},t\right)\right\|^{2}+\frac{1}{2\sigma^{2}}\left\|r\left(\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{t},t\right)\right)\right\|^{2}],$
where, the detailed proof of the following equation can be found in [ho2020denoising].
$E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]=E_{t\sim[1,T],\mathbf{x}_{0:T}\sim q\left(\mathbf{x}_{0:T}\right)}\left[\frac{\tilde{A}_{t}^{2}}{2\tilde{\beta}_{t}^{2}}\left\|\mathbf{x}_{0}-\mathbf{x}_{\theta}\left(\mathbf{x}_{t},t\right)\right\|^{2}\right].$